OpenStudy (anonymous):

write the equivalent measure of the given angle: 1.)45 degrees. 2) π 3) 7 π/ 6

OpenStudy (anonymous):

equivalent degree and radian measures

hartnn (hartnn):

\(\Large \pi^r = 180^o\)

OpenStudy (anonymous):

hii @hartnn ! ^_^

hartnn (hartnn):

so, Radians to degrees, \(\Large x^r = x\times 180^0/\pi\) Degrees to radians , \(\Large x^o = x \times (\pi/180)\) oh hai dara :)

hartnn (hartnn):

45 is in degrees so use 2nd formula, put x =45

OpenStudy (anonymous):

heyy, i am using this formula, radians to degrees : 180 degrees/ πrad

OpenStudy (anonymous):

degrees to radians : πrad / 180 degrees

OpenStudy (anonymous):

aww this is just the same Lol

hartnn (hartnn):

for radians to degrees , that what i wrote, 1st formula, multiply by 180, divide it by pi for degrees to radians .... yes same lol

OpenStudy (anonymous):

is my solution correct? given :π sol'n: πrad x 180 degrees / πrad = 180

hartnn (hartnn):

yes! pi radians = 180 degrees , correct :)

OpenStudy (anonymous):

45 degrees = 45 degrees x πrad/180 degrees = 45 π/180 rad = =π/4 rad

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

:) :) :) :) :) okayy

OpenStudy (anonymous):

and the 3rd problem?

hartnn (hartnn):

you try first ? multiply by 180, and divide it by pi ....

OpenStudy (anonymous):

i got it

hartnn (hartnn):

nice! :)

OpenStudy (anonymous):

270pi /180 rad = 3/2 rad?

hartnn (hartnn):

what was the question for that ?

hartnn (hartnn):

270 degrees ?

OpenStudy (anonymous):

for 270 degrees

hartnn (hartnn):

yes, then where did pi go in the answer :P

OpenStudy (anonymous):

3 π / 2 rad

hartnn (hartnn):

now correct :)

OpenStudy (anonymous):

is it necessary to put a rad?

hartnn (hartnn):

its the unit. like if you say i walked 2 , will it sound good ? i walked 2km is more accurate right ?

hartnn (hartnn):

so, yes, necessary to put rad

OpenStudy (anonymous):

okayy. thaanks

hartnn (hartnn):

wow , that took you 10 minutes to understand :P lol, just kidding ...

hartnn (hartnn):

what were you typing ? :P sorry i didn't mean to interrupt...

OpenStudy (anonymous):

how about determining the quadrants of the angle. the given is -146 degrees. and the other one is -328.

hartnn (hartnn):

you know the range of angles in every quadrant ?

OpenStudy (anonymous):

sorry, i'm chatting with my friend lately so it took m e almost 10 minutes. lol

OpenStudy (anonymous):

quadrant 1 if i

OpenStudy (anonymous):

t's below 90 deg?

hartnn (hartnn):